OpenStudy (anonymous):

(x^(3/7) ) / (x^(5/7) )

OpenStudy (anonymous):

This problem can be solved using the properties of exponents: (x^a)/(x^b)= x^(a-b) (x^(3/7))/(x^(5/7)= x^(3/7-5/7) = X^(-2/7)